Calculate the magnetic flux on a square sheet of slid 0.2m and on which a magnetic field of 2T is felt.
1
Expert's answer
2019-09-26T14:55:47-0400

According to the definition of magnetic flux "F = Ba^2 = 2\cdot 0.04 = 0.08 (Wb)."